OpenStudy (anonymous):

I WILL MEDAL! As a salesperson at Trading Cards Unlimited, Justin receives a monthly base pay plus commission on all that he sells. If he sells $400 worth of merchandise in one month he is paid $292. If he sells $900 worth of merchandise in one month he is paid $407. Find Justins salary when he sells $3,200 worth of merchandise.

OpenStudy (aihberkhan):

Okay. Let x = monthly base pay r = the commission as a decimal y = the amount paid monthly If he sells $400 worth of merchandise in one month, he is paid $292: \[292 = x + r*400\] If he sells $900 of merchandise in one month, he is paid $407: \[407 = x + r*900\] We have the following system of equations: \[x + r*400 = 292\] \[x + r*900 = 407\] \[x = $200\] r = 0.23 or 23% commission \[y = 200 + 0.23x\] When he sells $3,200 worth of merchandise we have x = $3,200 and we need to find y: \[y = 200 + 0.23*3200\] \[y = $936\] Justin's salary when he sells $3,200 worth of merchandise is $936.

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @skyyxlar

OpenStudy (anonymous):

Thankyou!! I got the answer! @AihberKhan

OpenStudy (aihberkhan):

No problem! @skyyxlar